OpenStudy (vera_ewing):

Which of the following is a solution of x2 - 10x = -36?

OpenStudy (vera_ewing):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

the quadratic formula is (-b +- sqrt(b^2-4ac))/2a

OpenStudy (anonymous):

either ( x - 8 ) = 0 or ( x - 2 )= 0

OpenStudy (alexandervonhumboldt2):

nope

OpenStudy (vera_ewing):

@TheSmartOne is it going to be +5 and +i or -5 and -i?

TheSmartOne (thesmartone):

Quadratic Formula is \(\Huge\sf \frac{-b \pm \sqrt{b^2-4ac}}{2a}\)

TheSmartOne (thesmartone):

\(\sf a=1\\ b=-10 \\c=36\)

TheSmartOne (thesmartone):

Plug in the numbers and we get \(\sf \LARGE\frac{ -(-10) \pm \sqrt{ (-10)^2 - 4 \cdot 1 \cdot 36} }{ 2 \cdot 1 }\)

TheSmartOne (thesmartone):

And just simply that and tell me what you get @vera_ewing